Jefferson Lab HALL A Logbook for December 04 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

    
    

    Wednesday

    
    

    Tuesday

    
    

    Wednesday

    
    

    Tuesday

    
    

    Monday

    
    

    Sunday

  • 138159 12/12/04 12:44 R. Michaels new caget and caput
  • Friday

  • 138158 12/10/04 11:10 R. Michaels duplicate DAQ files
  • Thursday

    
    

    Wednesday

    
    

    Tuesday

    
    

    Monday

  • 138036 12/06/04 08:02 fieschi shift summary
  • 138033 12/06/04 07:24 Carlos Target moved to empty
  • 138030 12/06/04 07:20 Carlos The end...
  • 138023 12/06/04 05:22 fieschi datamon and xscalers
  • 138022 12/06/04 05:20 fieschi DC current for neutron veto
  • 138021 12/06/04 05:19 fieschi DC current and HV for proton array
  • 138020 12/06/04 05:17 fieschi DC current and HV for calorimeter
  • 138019 12/06/04 05:17 fieschi hall A tools
  • 138006 12/06/04 00:07 Carlos Swing shift summary
  • Sunday

  • 138004 12/05/04 23:46 Carlos Compton snapshot
  • 138003 12/05/04 23:45 Carlos xscaler and datamon snapshot
  • 138002 12/05/04 23:44 Carlos PA/Tagger HV
  • 138001 12/05/04 23:43 Carlos Calorimeter HV and ADC snapshots
  • 138000 12/05/04 23:42 Carlos Hall A tools snapshot
  • 137999 12/05/04 23:39 fieschi DC current for neutron veto
  • 137998 12/05/04 23:38 fieschi HV and DC current for proton array
  • 137964 12/05/04 16:17 ru cryotarget status 2
  • 137963 12/05/04 16:17 ru cryotarget status 1
  • 137960 12/05/04 16:07 Eric Voutier Shift summary
  • 137957 12/05/04 15:22 lkaufman Proton array/tagger ADC snapshot
  • 137934 12/05/04 08:02 fieschi shift summary
  • 137925 12/05/04 05:29 Moffit Array/Tagger ADC snapshot
  • 137920 12/05/04 04:13 Moffit Target: Moved to 15cm LH2
  • 137908 12/05/04 00:22 Carlos Swing shift summary
  • Saturday

  • 137877 12/04/04 18:41 Carlos /adaql3/work1 almost full. Some files deleted.
  • 137869 12/04/04 17:26 ru cryotarget status 2
  • 137868 12/04/04 17:25 ru cryotarget status 1
  • 137858 12/04/04 16:48 Carlos Optics runs 4919 and 5033
  • 137855 12/04/04 16:01 Eric Voutier Shift summary
  • 137847 12/04/04 14:42 dhayes PA/tagger ADCs
  • 137831 12/04/04 10:57 R. Michaels auto abu repaired
  • 137823 12/04/04 08:04 fieschi shift summary
  • 137814 12/04/04 06:00 Moffit Array/Tagger ADC snapshot
  • 137806 12/04/04 04:33 R. Michaels/ 192 network switch replaced
  • 137800 12/04/04 01:39 R. Michaels halladaq-cat switch broken
  • 137799 12/04/04 00:15 Carlos Swing shift summary
  • 137798 12/04/04 00:14 A. Camsonne ER1 error message
  • Friday

  • 137783 12/03/04 22:57 Carlos ROC17 crashed
  • 137777 12/03/04 20:52 hyde Proton Array ADC
  • 137770 12/03/04 18:24 ru cryotarget status
  • 137769 12/03/04 18:22 ru cryotarget status
  • 137762 12/03/04 16:14 hyde Proton Array HV / ADC
  • 137752 12/03/04 15:58 Eric Voutier Shift Summary
  • 137746 12/03/04 14:34 jpchen back to 15 cm LD2 target
  • 137743 12/03/04 14:24 jpchen moved to optics target
  • 137742 12/03/04 14:23 Eugene Moller results
  • 137733 12/03/04 11:20 jpchen cryotarget status (II)
  • 137732 12/03/04 11:19 jpchen cryotarget status
  • 137724 12/03/04 08:05 fieschi Shift summary
  • 137716 12/03/04 05:31 Moffit Array/Tagger ADC snapshot
  • 137703 12/03/04 00:47 Carlos Swing shift summary
  • Thursday

  • 137692 12/02/04 21:00 H.S. Jo Proton Array / Tagger ADC snapshot
  • 137689 12/02/04 20:25 Carlos Golden runs updated
  • 137677 12/02/04 16:52 Carlos Target moved to LD2 15cm
  • 137675 12/02/04 16:46 Carlos HARP scan results
  • 137673 12/02/04 16:34 Carlos Left Q2 cycled
  • 137669 12/02/04 16:03 Eric Voutier Shift summary
  • 137668 12/02/04 15:58 lkaufman Target moved to optics
  • 137667 12/02/04 15:40 lkaufman Proton array/tagger snapshot
  • 137664 12/02/04 14:54 lkaufman Target fan speed and JT valve changes
  • 137663 12/02/04 12:51 lkaufman LHRS p0 changed to 2.95
  • 137662 12/02/04 11:42 lkaufman LHRS moved to 19.32 deg
  • 137661 12/02/04 10:44 Eric Voutier PA & Tagger HV
  • 137660 12/02/04 10:13 lkaufman Target moved to empty
  • 137658 12/02/04 09:58 Eric Voutier Proton array DC snapshot
  • 137657 12/02/04 09:55 Eric Voutier Calorimeter DC snapshot
  • 137656 12/02/04 09:53 A. Camsonne Scalers at 2.5 uA Deuterium kin#3
  • 137655 12/02/04 09:52 Eric Voutier Tagger DC snapshot
  • 137653 12/02/04 09:37 lkaufman Target moved to LD2
  • 137649 12/02/04 08:01 fieschi shift summary
  • 137641 12/02/04 05:20 Moffit Array/Tagger ADC snapshot
  • 137632 12/02/04 00:05 A. Deur Swing shift summary
  • Wednesday

  • 137623 12/01/04 22:55 H.S. Jo Shape of the Proton Array channel 53
  • 137620 12/01/04 21:58 A. Deur /adaql3/work1/adaq/onlroot is full
  • 137611 12/01/04 18:00 Carlos Golden runs updated
  • 137610 12/01/04 17:34 A. Camsonne adaql2 filesystem problem
  • 137609 12/01/04 17:32 R. Michaels adaql2 rebooted to clean "/"
  • 137607 12/01/04 17:03 A. Deur Proton array and tagger ADC
  • 137606 12/01/04 16:02 mazouz Shift summary
  • 137604 12/01/04 15:43 mazouz TS1 crahed
  • 137603 12/01/04 13:08 Carlos Calorimeter HV change
  • 137600 12/01/04 12:37 mazouz no signals for B26 (tagger)
  • 137596 12/01/04 10:57 mazouz PA DC current & HV
  • 137595 12/01/04 10:57 mazouz Calo DC current & HV
  • 137594 12/01/04 10:56 mazouz 2.5 uA beam
  • 137593 12/01/04 10:35 mazouz Tagger DC current at 2uA and 4 uA
  • 137591 12/01/04 10:14 lkaufman Target moved to15 cm Hydrogen
  • 137590 12/01/04 10:12 lkaufman LHRS p0 changed to 2.35 Gev/c
  • 137589 12/01/04 08:15 Carlos Owl shift summary
  • 137585 12/01/04 06:45 Moffit Target to Carbon..
  • 137583 12/01/04 06:36 Moffit Array/Tagger ADC snapshot
  • 137581 12/01/04 03:54 Moffit Target to "Optics"
  • 137578 12/01/04 02:31 Moffit Left HRS p0 setting
  • 137576 12/01/04 01:19 Carlos Compton vertical scan
  • 137573 12/01/04 00:08 fieschi shift summary